### v1

# Rethinking CLIP-based Video Learners in Cross-Domain Open-Vocabulary Action Recognition

Abstract
Contrastive Language-Image Pretraining (CLIP) has shown remarkable open-vocabulary abilities across various image understanding tasks. Building upon this impressive success, recent pioneer works have proposed to adapt the powerful CLIP to video data, leading to efficient and effective video learners for open-vocabulary action recognition. Inspired by the fact that humans perform actions in diverse environments, our work delves into an intriguing question: Can CLIP-based video learners effectively generalize to video domains they have not encountered during training? To answer this, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, and conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners under various types of domain gaps. Our evaluation demonstrates that previous methods exhibit limited action recognition performance in unseen video domains, revealing potential challenges of the cross-domain open-vocabulary action recognition task. To address this task, our work focuses on a critical challenge, namely scene bias, and we accordingly contribute a novel scene-aware video-text alignment method. Our key idea is to distinguish video representations apart from scene-encoded text representations, aiming to learn scene-agnostic video representations for recognizing actions across domains. Extensive experimental results demonstrate the effectiveness of our method. The benchmark and code will be available at https://github.com/KunyuLin/XOV-Action/.
1 Introduction
Figure 1: (a) Can CLIP-based video learners effectively generalize to unseen test domains? For example, given a model trained with normal videos, we wonder if it can effectively recognize actions in dark videos. (b) We conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners on four test datasets, namely UCF [59], HMDB [30], ARID [78] and NEC-Dr [12]. For each test dataset, we report the accuracy of closed-set and open-set action categories according to the training categories in Kinetics400 [6]. As shown above, these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains. Note that, for each metric, we report the best performance among all methods (namely different checkpoints), please refer to Table 3 for the full results. Best viewed in color.
Action recognition aims to recognize what actions humans are performing in videos, which has wide applications in surveillance systems, health monitoring, sport analysis, etc [28, 62, 14, 13, 94, 17, 22]. Recently, inspired by the impressive success of Contrastive Language-Image Pretraining (CLIP) [52] across various image understanding tasks, pioneer works propose to adapt CLIP to video data for action recognition [72, 46, 26, 75, 53, 73]. Different from traditional models that focus on closed-set recognition [56, 71, 38, 6, 3], the new CLIP-based recognition paradigm leads to efficient video learners with remarkable open-vocabulary recognition abilities, i.e., they achieve state-of-the-art performance for both closed-set and open-set action categories on various popular video datasets with moderate training cost. Such open-vocabulary abilities significantly enhance the practical relevance of action recognition models.
As shown in Figure 1(a), in this work, we delve into an intriguing question: Can CLIP-based video learners effectively generalize to video domains that they have not encountered during training? The generalization ability in unseen video domains is crucial for action recognition models in practice, since models often suffer from environment changes when deployed in real-world applications [7, 85]. For example, surveillance systems will encounter actions performed under illumination shifts caused by day-night change or weather change. Therefore, from a practical perspective, we expect that a CLIP-based video learner possessing open-vocabulary abilities can robustly adapt to domain changes for action recognition.
To answer this, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, where the four test datasets exhibit various levels of domain gaps in comparison to the training datasets. Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners for cross-domain open-vocabulary action recognition. As shown in Figure 1(b), our evaluation reveals that these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains. Specifically, these models exhibit very limited performance for closed-set categories in test domains with large domain gap (e.g., ACC of 53.89% for dark videos in ARID). Also, we note that the open-set recognition performance of these CLIP-based video learners is far from reaching saturation even in domains with moderate domain gap (i.e., ACC of 42.22% on HMDB). Overall, cross-domain open-vocabulary action recognition is challenging and underexplored.
To address this task, our work focuses on tackling the domain gaps by mitigating scene bias, which is a critical challenge of the task. Specifically, the scene bias is caused by the high correlation between human actions and specific scenes in training videos, also known as the spurious correlation [36, 11, 69, 33]. For example, as shown in Figure 1(a), humans usually perform jumping in track and field in the training domain (Kinetics400), thus models prefer to recognize the action according to the track and field since the static scenes are easier to fit [36, 11, 69]. However, humans may perform jumping in hallways (ARID) or basketball courts (NEC-Dr), thus recognizing “jump” by the track and field would result in recognition errors across domains.
To mitigate the scene bias, we propose a novel Scene-Aware video-Text Alignment (SATA) method for learning scene-agnostic video representations. Specifically, we propose to distinguish video representations apart from scene-encoded text representations, which encourages the video encoder to downweight the attention on scene information and thus pay more attention to action information.
In summary, our contributions are listed as follows:
  • We establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, including four test domains exhibiting various levels of domain gap with respect to the training domains. We identify closed-set and open-set categories for each test domain, thus providing a comprehensive way to evaluate models across various situations.
  • We conduct a thorough evaluation of five state-of-the-art CLIP-based video learners based on XOV-Action. Our experimental analysis demonstrates the limited performance of these methods in cross-domain open-vocabulary action recognition, revealing potential challenges of this task.
  • We propose a novel scene-aware video-text alignment method, which learns scene-agnostic video representations to recognize actions across domains. And, extensive experiments demonstrate the effectiveness of our method.
2 Related Works
2.1 Action Recognition
Action recognition aims to recognize human actions in videos. In the past decade, motivated by the success of deep learning [29, 57, 64, 21, 63, 10, 68], many video classification architectures have been proposed. Among these architectures, there are three most popular types, namely 2D CNNs, 3D CNNs and Video Transformers. Typically, 2D CNNs adopt 2D convolution for spatial modeling, and conduct temporal modeling beyond spatial modeling [71, 93] or embed temporal shift into spatial modeling [38, 55, 60]. Different from 2D models, 3D CNNs extends 2D convolution to 3D convolution for adapting video data [65, 6, 66, 67, 16, 90, 35]. By adopting attention mechanisms, Video Transformers expand the receptive field of 2D and 3D CNNs, leading to remarkable performance [19, 45, 89, 91, 3, 2]. Although above models show promising performance for closed-set action recognition, they usually lack the ability to recognize open-set categories.
2.2 Cross-Domain Action Recognition
Cross-domain action recognition aims to learning video classification models by transferring knowledge from source domains to target domains. This area includes two tasks, namely domain-adaptive and domain-generalizable action recognition. In domain-adaptive action recognition [61, 77, 25, 7, 79, 5, 80], unlabeled videos from target domains are accessible for training, and thus related works usually focusing on developing models oriented to specific target domains. Typical works address domain-adaptive action recognition by learning cross-domain invariance, i.e., exploring the shared action cues between the source and target domains [7, 48, 48, 81, 42, 54]. Other works usually focus on leveraging the multi-modal nature of video data [44, 58, 27, 84, 92].
Different from domain-adaptive action recognition, domain-generalizable action recognition [85, 50, 39, 51] aims to learn generalizable models in unseen test domains, i.e., videos of the target domains are not accessible during training. In the absence of target videos, existing works usually assume specific types of invariance across domains for generalizable action representation learning, e.g., Yao et al. assume that local temporal features are more invariant across domains compared with global temporal features [85]. Instead of invariance mining, Lin et al. propose to learn diverse action features for generalization in unseen target domains [39]. Our cross-domain open-vocabulary action recognition task is closely related to domain-generalizable action recognition, since test domains are unseen during training. However, our task is much more challenging as we strive for open-vocabulary abilities, leading to significantly different technical designs.
2.3 Vision-Language Pretraining
Vision-language pretraining [52, 34, 37, 1, 88, 86, 15, 83, 18] has made great progress in recent years, and one of the most impressive works is Contrastive Language-Image Pretraining (CLIP) [52]. By utilizing web-scale paired image-text data for training, CLIP shows robust zero-shot object recognition abilities. Also, some advanced works have demonstrated that CLIP can effectively solve specific downstream tasks by efficient adaptation [97, 87, 49, 96, 53, 41]. Moreover, integrating CLIP with specialized techniques shows remarkable open-vocabulary abilities on various image understanding tasks, such as object detection and segmentation [52, 74, 20, 31, 40]. Although large-scale image-text pretraining has achieved great success, video-text pretraining still has room for development [76, 32, 70, 43, 23, 9, 53]. It is because videos are inherently more complex than images, and large-scale paired video-text datasets are less available. Therefore, it is valuable to develop efficient methods for adapting pretrained image-text models to video understanding tasks.
3 The Proposed XOV-Action Benchmark
Our work focuses on the cross-domain open-vocabulary action recognition task. This task aims to develop open-vocabulary action recognition models that are generalizable in unseen target domains by training in the source domain, i.e., recognizing both closed-set and open-set action categories in new test domains. This task, closely aligning with real-world applications, requires action recognition models to exhibit open-vocabulary classification abilities and demonstrate robustness in handling environmental changes. To evaluate models on this task, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, which provides a comprehensive way to analyze models across various situations. In what follows, we introduce the components of our XOV-Action benchmark in detail, as well as evaluation metrics and potential challenges.
Benchmark Components: Our proposed XOV-Action benchmark consists of two source datasets for training and four target datasets for test. Specifically, the two source datasets used for training are as follows:
(1) Kinetics400 [6]: One of the most widely-used action recognition datasets, consisting of 400 action categories. Videos in Kinetics400 are collected from YouTube, which are usually recorded in normal environments (e.g., normal illumination and weather). Existing works of CLIP-based video learners typically use Kinetics400 for training, and we use the original train-validation split following them.
(2) Kinetics150: A subset of Kinetics400, composed of 150 action categories selected from the full Kinetics400. These 150 categories include all the closed-set categories of the four target datasets in comparison to Kinetics400, and the remaining categories are random sampled. We will illustrate the definition of closed-set categories later. We construct this subset for a detailed analysis to cross-domain open-vocabulary action recognition models.
The four target datasets for test are as follows:
(1) UCF [59]: One of the most widely-used action recognition datasets, consisting of 101 action categories. Videos in UCF are collected from YouTube, and videos of each action are usually captured from specific or similar environments. UCF has a small domain gap compared with the Kinetics (source) domain, and previous CLIP-based video learners commonly use UCF to evaluate their open-vocabulary recognition abilities.
(2) HMDB [30]: A widely-used action recognition datasets consisting of 51 action categories. Compared with UCF, videos in HMDB are captured from more unconstrained environments and more different camera views. Specifically, Videos in HMDB are collected mainly from movies, and remaining videos are from Prelinger archive, YouTube or Google videos. Overall, HMDB has a moderate domain gap compared with the Kinetics domain [80, 82].
(3) ARID [78]: A dataset consisting of 11 categories of action videos, which are recorded under dark environments. These actions include singular person actions (e.g., jump, run) and actions associated with objects (e.g., drink, pick). Due to the significant different illumination conditions, ARID exhibits a large domain gap compared with the Kinetics domain (see Figure 1 for an example). Developing models that are generalizable in ARID is challenging, when training data are limited to videos with normal illumination.
(4) NEC-Dr [12]: A dataset consisting of 16 categories of action videos, which are recorded by drones in the same basketball court. These actions include single-person actions (e.g., jump, walk) and interactive actions (e.g., hug, shake hands). Due to the different shooting equipments and scenarios, NEC-Dr exhibits a large domain gap compared with the Kinetics domain (see Figure 1 for an example). Therefore, it is challenging to make a Kinetics-trained model generalizable in such a drone video domain.
In our experiments, we use videos from Kinetics400 or Kinetics150 for training. Then, we evaluate models on the four target datasets, which exhibit various levels of domain gap in comparison to the training datasets. Please refer to the Appendix for more details of our XOV-Action.
Evaluation Metrics: First, we illustrate the definitions of closed-set and open-set categories for test domains: (a) The closed-set categories refer to the categories that share similar meanings and have a common lexicon with the categories in the training domain, following a similar approach in previous zero-shot action recognition works [4, 8, 95]. (b) The open-set categories refer to the remaining categories that are not involved in the training domain. Then, we identify closed-set and open-set categories for each test domain, according to the Kinetics400 categories. Category statistics are summarized in Table 1.
In our experiments, we adopt three evaluation metrics:
(1) The closed-set accuracy measures the recognition performance of closed-set categories, which primarily evaluates the model abilities of tackling domain gaps.
(2) The open-set accuracy measures the performance of open-set categories, which evaluates the generalization abilities across both video domains and action categories.
(3) The overall accuracy measures the recognition performance over all categories, which provides a holistic view of model effectiveness across various situations.
Note that, some previous works of CLIP-based video learners treat all the categories in UCF and HMDB as open-set for models trained on Kinetics400. However, UCF and HMDB have many overlapping categories with Kinetics400, thus these works make an inaccurate assessment of the model’s open-vocabulary abilities. Unlike these works, we provide a more accurate way for evaluation by distinguishing between closed-set and open-set categories. By evaluating across various domain gaps and different action categories, our proposed XOV-Action benchmark enables us to perform a wide range of analysis for cross-domain open-vocabulary action recognition models.
Potential Challenges: Although cross-domain open-vocabulary action recognition is valuable and practical, it is challenging and underexplored. Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners, as shown in Table 3. The experimental results demonstrate that these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains. Specifically, the results reveal three fundamental challenges of the cross-domain open-vocabulary action recognition task as follows:
(1) How to tackle the domain gap between the training and test domains? Previous CLIP-based video learners exhibit very limited performance in test domains with large domain gaps, i.e., the best closed-set ACCs are only 53.89% and 31.48% on ARID and NEC-Dr, respectively.
(2) How to improve the generalization abilities for open-set categories in the presence of domain gaps? The cross-domain open-set recognition performance is far from reaching saturation even in test domains with moderate domain gaps, e.g., the best open-set ACC is only 42.22% on HMDB.
(3) How to develop a single model that achieves superior performance on all the test domains? It is very challenging since existing models demonstrate varied performance across various test domains as shown in Table 3.
Table 1: Category statistics of four test domains. Closed-set and open-set categories are identified according to Kinetics400.
Test Domains UCF HMDB ARID NEC-Dr
# of closed-set actions 50 33 6 7
# of open-set actions 51 18 5 9
# of all actions 101 51 11 16
Domain gap Small Moderate Large Large
4 Methodology
In this work, we focus on tackling the domain gaps for cross-domain open-vocabulary action recognition. Specifically, we propose to address a critical challenge of the task, namely scene bias. To this end, we propose a novel Scene-Aware video-Text Alignment (SATA) method. In what follows, we illustrate our method in detail.
Figure 2: An overview of our proposed Scene-Aware video-Text Alignment (SATA) method. Our method basically includes a video encoder and a text encoder for video and text representation extraction, with a contrastive loss for video-text alignment. Based on the scene-encoded text prompts, we propose the Scene-Aware Discrimination and Scene-Aware Consistency losses, aiming to learn scene-agnostic video representations for cross-domain open-vocabulary action recognition. Best viewed in color.
4.1 Problem Formulation
In cross-domain open-vocabulary action recognition, a set of labeled videos $\mathcal{D}=\{(x,y)\}$ from a source domain are given for training, where $x$ and $y$ denote a source video and its ground-truth label index. The source domain consists of $K$ action categories, and the action names (label texts) are denoted by $\{a_{1},a_{2},\dots,a_{K}\}$. Given only source videos for training, our goal is to develop a model that is generalizable in unseen target domains, where the source and target domains follow different distributions and different label spaces. Following previous open-vocabulary action recognition works, we sample $T$ frames from each video as model input during training and test, i.e., $x=\{x_{1},x_{2},\dots,x_{T}\}$.
4.2 Model Overview
Building upon CLIP, our model consists of a video encoder $f_{\text{vid}}(\cdot)$ and a text encoder $f_{\text{txt}}(\cdot)$. Following OpenVCLIP [73], we construct a video encoder by integrating patches from neighboring frames in each self-attention layer of the original CLIP image encoder. For video $x$, the video representation $e_{x}$ is obtained by the average of local video representations, i.e., $e_{x}=f_{\text{vid}}(x)=1/T\sum_{t=1}^{T}f^{t}_{\text{vid}}(x)$, where each local representation corresponds to one frame. We use the original CLIP text encoder and keep it frozen during training. Following previous works [73, 53], we use a video-text alignment loss to adapt video data, which is formulated as follows:
$L_{\text{vta}}=-\log{\frac{\exp{(s(e_{x},e_{a_{y}})/\tau)}}{\sum_{k=1}^{K}\exp{(s(e_{x},e_{a_{k}})/\tau)}}},$ (1)
where $s(\cdot,\cdot)$ denotes the cosine similarity, $\tau$ is the temperature, $e_{a_{k}}=f_{\text{txt}}(g(a_{k}))$ is the text representation of the $k$-th action category. The function $g(\cdot)$ transforms an action name into a text prompt in the form of “a video of a person [doing something].”, e.g., “a video of a person abseiling.” for the action “abseiling”. To mitigate scene bias, we contribute a Scene-Aware Discrimination loss and a Scene-Aware Consistency loss for learning scene-agnostic video representations, which will be illustrated in the next subsection. An overview of our proposed SATA is shown in Figure 2.
4.3 Scene-Aware Video-Text Alignment
To learn scene-agnostic video representations for cross-domain open-vocabulary action recognition, we contribute a novel Scene-Aware video-Text Alignment (SATA) method. The key idea of our SATA is to distinguish video representations apart from scene-encoded text representations, which encourages the video encoder to downweight the attention on scene information in videos.
Scene-Aware Discrimination: First of all, we randomly sample $N$ scene suffixes and construct scene-encoded text prompts for each training video. Each scene suffix is in the form of “[at/on/in the/a scene]”, e.g., “in the park”, “on the street”. In our implementation, we ask GPT-4 [47] to automatically generate a pool of scene suffixes for random sampling, which involves no human annotation cost. Based on these suffixes, we construct scene-encoded text prompts of ground-truth action category for each video, which is in the form of “a video of a person [doing something] [at/on/in the/a scene].” For example, for the action “abseiling”, we obtain $N$ scene-encoded text prompts, e.g., “a video of a person abseiling in the park.”.
Table 2: Evaluation of CLIP-based video learners trained on Kinetics150 on our XOV-Action benchmark. “C”, “O” and “A” denote the accuracy (%) of closed-set, open-set and all action categories, respectively. “AVG” denotes the average performance over the four test domains. * indicates the model variant with frozen text encoder. The bold and underlined numbers indicate the best and second best.
Models UCF HMDB ARID NEC-Dr AVG
C O A C O A C O A C O A C O A
CLIP 61.51 57.85 59.67 42.92 38.62 41.40 32.04 12.04 24.02 18.52 6.44 12.33 38.75 28.74 34.35
   ActionCLIP [72] 84.76 56.27 70.44 56.78 33.02 48.40 42.25 19.75 33.23 18.75 4.22 11.30 50.63 28.32 40.84
LSTM [72] 84.54 58.38 71.39 58.30 36.38 50.56 42.04 26.85 35.95 17.13 12.44 14.73 50.50 33.51 43.16
TConv [72] 84.54 58.61 71.24 57.29 36.19 49.84 44.29 27.78 37.67 17.59 12.22 14.84 50.93 33.56 43.40
XCLIP [46] 89.89 59.38 74.59 53.44 30.78 45.62 35.10 13.72 26.64 16.90 5.70 11.26 48.83 27.40 39.53
Text4Vis [75] 83.40 49.84 66.54 54.25 38.06 48.54 41.43 31.71 37.53 12.50 4.17 8.23 47.90 30.94 40.21
ViFiCLIP [53] 87.77 49.32 68.44 61.94 27.05 49.63 40.20 21.95 32.89 14.58 0.00 7.10 51.12 24.58 39.52
ViFiCLIP [53] 82.93 62.07 72.44 58.60 41.60 52.60 44.69 27.13 37.65 17.36 10.53 13.86 50.90 35.33 44.14
OpenVCLIP [73] 82.98 63.17 73.03 56.48 44.96 52.41 48.16 27.13 39.73 19.68 13.38 16.45 51.82 37.16 45.40
SATA (Ours) 85.53 62.01 73.71 58.70 44.78 53.79 47.35 33.54 41.81 27.55 10.97 19.04 54.78 37.82 47.09
Then, based on the scene-encoded text prompts, we design a Scene-Aware Discrimination loss, which is given as follows:
$L_{\text{dis}}=-\log{\frac{\exp{(s(e_{x},e_{a_{y}}))}}{\exp{(s(e_{x},e_{a_{y}}))}+\sum_{n=1}^{N}\exp{(s(e_{x},\tilde{e}^{n}_{a_{y}}))}}}.$ (2)
In this loss, $\tilde{e}^{n}_{a_{y}}=f_{\text{txt}}(\tilde{g}(a_{y},n))$ is the representation of a scene-encoded text prompt, which encodes semantic information of both the ground-truth action $a_{y}$ and a random scene. The function $\tilde{g}(\cdot,n)$ transforms an action name into a scene-encoded text prompt using the $n$-th scene suffix. Accordingly, our Scene-Aware Discrimination loss pushes the video representations away from the scene-encoded text representations in video-text alignment. In this way, the loss leverages the strong power of CLIP text encoder, which encourages the video encoder to pay less attention on scene information and thus pay more attention to action information, thereby bridging the domain gap.
Scene-Aware Consistency: In addition, we propose a Scene-Aware Consistency loss to constrain the video representation learning. Our intuition is that a video should have a more consistent representation with its scene-encoded text prompts compared with text representations of other action categories, since the scene-encoded text prompts encodes semantic information of the ground-truth action category (in addition to scenes). For example, consider a training video that depicts a person playing basketball in a court. The video should have a similar representation to the scene-encoded text prompts like “a video of a person playing basketball in the park.”, compared with text prompts of other action categories, e.g., “a video of a person playing football.”. Specifically, our proposed Scene-Aware Consistency loss is formulated as follows:
$L_{\text{con}}=$ $-\frac{1}{N+1}\log{\frac{\exp{(s(\hat{e}_{x},e_{a_{y}}))}}{Z}}$ (3)
$-\frac{1}{N+1}\sum\nolimits_{n=1}^{N}\log{\frac{\exp{(s(\hat{e}_{x},\tilde{e}^{n}_{a_{y}})))}}{Z}},$
where $\hat{e}_{x}=f_{\text{proj}}(f_{\text{vid}}(x))$ is the video representation after projection $f_{\text{proj}}(\cdot)$, and $Z=\sum_{k=1}^{K}\exp{(s(\hat{e}_{x},e_{a_{k}}))}+\sum_{n=1}^{N}\exp{(s(\hat{e}_{x},\tilde{e}^{n}_{a_{y}}))}$. Note that, the projection function is crucial to harmonize our Scene-Aware Discrimination and Scene-Aware Consistency losses, as shown in our experimental analysis.
Furthermore, we propose a Text-Adaptive Aggregation module to enhance video representations for classification. Specifically, we aggregate local video representations by text-adaptive weights instead of average pooling. The aggregation module is formulated as $e^{\text{txt}}_{x}=\sum_{t=1}^{T}w^{t}_{x}e^{t}_{x}$, where $e^{t}_{x}=f^{t}_{\text{vid}}(x)$ is the $t$-th local representation. And, the weight $w^{t}_{x}$ is as follows:
$w^{t}_{x}=\frac{1}{K}\sum_{k=1}^{K}\frac{\exp{(s(e^{t}_{x},e_{a_{k}})/\eta)}}{\sum_{m=1}^{T}\exp{(s(e^{m}_{x},e_{a_{k}})/\eta)}},$ (4)
where $\eta$ is the temperature. In this way, we obtain integrated video representations adaptive to specific action categories. In our implementation, we use $e^{\text{txt}}_{x}$ instead of $e_{x}$ in Eq. (1) and Eq. (2) for training, and also use $e^{\text{txt}}_{x}$ for test.
The overall training loss of our SATA is given as follows:
$L=L_{\text{vta}}+\lambda_{\text{dis}}L_{\text{dis}}+\lambda_{\text{con}}L_{\text{con}},$ (5)
where $\lambda_{\text{dis}}$ and $\lambda_{\text{con}}$ are coefficients for trade-off. During test, we use text presentations of action categories in target domains for classification.
Table 3: Evaluation of CLIP-based video learners trained on Kinetics400 on our XOV-Action benchmark. We use the officially released checkpoints of previous methods for evaluation, and we use one temporal clip during inference for a fair comparison. “C”, “O” and “A” denote the ACC (%) of closed-set, open-set and all action categories, respectively. “AVG” denotes the average performance over the four test domains. * indicates the model variant with frozen text encoder. The bold and underlined numbers indicate the best and second best.
Models UCF HMDB ARID NEC-Dr AVG
C O A C O A C O A C O A C O A
CLIP 66.71 59.53 63.15 49.19 35.26 44.27 37.76 12.65 27.69 10.42 22.44 16.59 41.04 32.47 37.92
   ActionCLIP [72] 91.13 59.76 75.36 64.24 27.78 51.37 48.68 29.09 40.83 27.88 15.54 21.55 57.98 33.04 47.28
XCLIP [46] 92.02 55.11 73.49 54.76 26.31 44.75 34.49 28.35 31.99 28.93 5.70 17.16 52.55 28.87 41.85
Text4Vis [75] 90.47 61.03 75.69 58.47 29.60 48.37 53.43 38.10 46.97 29.77 9.05 19.20 58.04 34.45 47.56
ViFiCLIP [53] 84.01 68.28 72.54 64.68 22.02 49.48 53.89 28.09 43.51 31.48 14.00 22.57 58.52 33.10 47.03
ViFiCLIP [53] 91.58 62.01 76.75 60.02 29.85 49.15 33.88 28.09 31.74 16.89 14.35 15.58 50.59 33.58 43.31
OpenVCLIP [73] 89.80 77.28 83.51 58.08 42.22 52.48 45.64 26.36 37.91 8.29 10.28 9.32 50.45 39.04 45.81
SATA (Ours) 88.51 65.17 76.78 61.44 44.03 55.29 54.49 28.05 43.89 28.70 11.84 20.06 58.29 37.27 49.01
Table 4: Ablation study of our proposed SATA trained on Kinetics150. “C”, “O” and “A” denotes the accuracy (%) of closed-set, open-set and all action categories, respectively. “AVG” denotes the average performance over the four test domains. The closed-set ACC primarily evaluates the model abilities of tackling domain gaps.
Models UCF HMDB ARID NEC-Dr AVG
C O A C O A C O A C O A C O A
Baseline 84.04 62.22 73.08 57.59 44.59 53.00 47.35 26.52 39.00 21.30 14.04 17.57 52.57 36.84 45.66
+$L_{\text{dis}}$ 84.63 60.43 72.47 59.41 43.10 53.65 49.18 25.31 39.61 25.93 10.53 18.03 54.79 34.84 45.94
+$L_{\text{dis}}$&$L_{\text{con}}$ 84.84 62.28 73.50 58.30 44.03 53.26 49.39 29.57 41.44 26.39 11.40 18.70 54.73 36.82 46.73
Full 85.53 62.01 73.71 58.70 44.78 53.79 47.35 33.54 41.81 27.55 10.97 19.04 54.78 37.82 47.09
+$L_{\text{dis}}$&$L_{\text{con}}$ w/o Proj 84.20 61.75 72.92 58.50 42.91 53.00 48.78 28.35 40.59 20.14 14.04 17.01 52.91 36.76 45.88
5 Experiments
5.1 Experimental Setups
Implementation details: For each video, our model takes $T=16$ frames of size $224\times 224$ as inputs. Following ViFiCLIP [53], we adopt temporal jitter, multi-scale random spatial crop and color jitter for augmentation during training. During test, we use one temporal clip composed of center-cropped frames for inference. Regarding the architecture of the video encoder, we adopt ViT-B/32 and ViT-B/16 initialized by CLIP image encoder for Kinetics150 and Kinetics400, respectively. For temporal modeling, we set the temporal receptive field of self-attention layers to 7 in the video encoder. We ask GPT-4 [47] to output 300 scene suffixes that are common in daily life, and randomly sample $N=50$ scene suffixes for Eq. (2) in each batch. The projection function $f_{\text{proj}}(\cdot)$ is an MLP with residual. The parameters are optimized using AdamW optimizer with a batch size of 512, a learning rate of 8e-6, a momentum of 0.9 and a weight decay of 1e-3. By default, we set the loss coefficients as $\lambda_{\text{dis}}=0.2$ and $\lambda_{\text{con}}=0.2$, and the temperature coefficients as $\tau=0.01$ and $\eta=0.1$. Following OpenVCLIP [73], we adopt the SWA technique [24] for producing the final model.
Evaluation protocol: We adopt five state-of-the-art CLIP-based video learners for evaluation and comparison. To conduct a solid evaluation on Kinetics400, we utilize the officially released checkpoints of each method, instead of retraining models. In the case of Kinetics150, we ensure a fair comparison by adopting consistent network architectures and training recipes across all the methods based on their official codes. During inference, we use one temporal clip for a fair comparison. We use our proposed XOV-Action benchmark for evaluation, which includes test domains of various types of domain gaps and action categories of different types. We report the accuracy of closed-set, open-set and all categories in individual test domains as well as the average accuracies, for a comprehensive analysis. When conducting a comparison between models, we prioritize the average ovarall accuracy as the primary metric.
5.2 Results
Comparison with state-of-the-arts: The results are summarized in Table 2 and 3. As shown in the tables, by adapting CLIP to video data, previous CLIP-based video learners obtain substantial improvement over the original CLIP for cross-domain open-vocabulary action recognition. Although these methods show impressive performance for the domain with a small gap (UCF), they still exhibit limited performance for the domains with large gaps (ARID and NEC-Dr). These results reveal potential challenges of the task, as discussed in Section 1 and 3. By learning scene-agnostic video representations, our proposed SATA outperforms previous state-of-the-arts by 1.69% and 1.45% in terms of average overall accuracy for Kinetics150 and Kinetics400, respectively. These improvements are significant considering that the metric represents an average across four accuracy metrics.
Figure 3: Quantitative analysis of the total number of scene suffixes.
Figure 3: Quantitative analysis of the total number of scene suffixes.
Figure 4: Quantitative analysis of the coefficient $\lambda_{\text{dis}}$ for Scene-Aware Discrimination loss by closed-set accuracy on four test domains. The horizontal axis shows the value of $\lambda_{\text{dis}}$.
Ablation Study: Table 4 summarizes the results of our ablation study. In the table, the baseline refers to the model trained with the basic video-text alignment loss $L_{\text{vta}}$. By introducing our proposed Scene-Aware Discrimination loss $L_{\text{dis}}$, our model obtains improvement on XOV-Action in terms of closed-set ACC on all the four test domains. Specifically, our model obtains significant improvements of 1.83% and 4.63% in terms of closed-set ACC on ARID and NEC-Dr, which have large domain gaps with the training domain. This is because that our proposed Scene-Aware Discrimination loss encourages the video encoder to downweight the attention on scene information in videos and thus pay more attention to action information that are generalizable across different video domains. However, we find that introducing only the Scene-Aware Discrimination loss reduces the open-set performance. By using both Scene-Aware Discrimination loss $L_{\text{dis}}$ and Scene-Aware Consistency loss $L_{\text{con}}$, our model can obtain improved closed-set performance while maintaining the open-set performance, and this leads to an improvement of 1.07% in terms of average overall accuracy over the baseline. Notably, the projection within our Scene-Aware Consistency loss is crucial to harmonize the two losses. If the projection is dropped, the improvement is very limited as shown in the last line of Table 4. Finally, by introducing the Text-Adaptive Aggregation module, our model obtains further improvement in terms of average overall accuracy (please refer to the Appendix for more analysis of this module).
Effect of scene suffixes: In Figure 4, we show how the suffix number affects the model performance. We set the loss coefficient $\lambda_{\text{dis}}=0.5$ in this experiment. As shown in the figure, our model obtains better closed-set accuracy as the total number of scene suffixes increases. According to the design of our Scene-Aware Discrimination loss, we distinguish videos apart from more scene-encoded text prompts in feature space with more scene suffixes used. In this way, the video representations are less likely to be confused with scenes, leading to stronger abilities for recognizing actions in unseen domains.
Figure 5: Analysis of the coefficient $\lambda_{\text{con}}$ for Scene-Aware Consistency loss. The horizontal axis shows the value of $\lambda_{\text{con}}$.
Effect of the Scene-Aware Discrimination loss: We conduct a quantitative analysis of the loss coefficient $\lambda_{\text{dis}}$, and the results on the four test domains are shown in Figure 4. The results in the figure show that our model consistently obtains better closed-set performance on all test domains with a larger loss weight. The comprehensive results demonstrate that our proposed Scene-Aware Discrimination loss effectively bridges the domain gaps of various types by mitigating scene bias.
Effect of the Scene-Aware Consistency loss: We conduct a quantitative analysis of the loss coefficient $\lambda_{\text{con}}$, and the results are shown in Figure 5. As shown in the figure, our model obtains higher open-set performance but lower closed-set performance when using a larger loss weight. The results empirically reveal that a good trade-off between our Scene-Aware Discrimination and Consistency losses is crucial for cross-domain open-vocabulary action recognition.
Figure 6: Qualitative analysis by attention visualization. Best viewed in color.
Qualitative analysis: In Figure 6, we show a qualitative analysis by attention visualization. As shown in the figure, compared with the baseline that pays much attention to scenes, our model focuses on the action performers when recognizing actions across domains. The results demonstrate that our model effectively mitigates scene bias.
5.3 Conclusion
This work concentrates on a valuable but underexplored task, namely cross-domain open-vocabulary action recognition. By contributing the XOV-Action benchmark, we provided a comprehensive way to evaluate models across various types of domain gaps and different action categories, and our evaluation of state-of-the-art CLIP-based video learners revealed the challenging nature of this valuable task. To address this task, we proposed a novel scene-aware video-text alignment method, aiming to tackle the domain gap by learning scene-agnostic video representations. Extensive experimental results demonstrated the effective of our method.
The cross-domain open-vocabulary action recognition task presents numerous challenges that remain to be addressed, such as cross-category generalization in cross-domain settings. Despite our efforts, we focus only on resolving scene bias in our current methodology. However, our proposed benchmark provides a comprehensive way for evaluation. We believe that our work will serve as a catalyst for further advancement, and we hope it will inspire future innovative solutions for this field.

### v2

@@
 Abstract
-Contrastive Language-Image Pretraining (CLIP) has shown remarkable open-vocabulary abilities across various image understanding tasks.
-Building upon this impressive success, recent pioneer works have proposed to adapt the powerful CLIP to video data, leading to efficient and effective video learners for open-vocabulary action recognition.
-Inspired by the fact that humans perform actions in diverse environments, our work delves into an intriguing question: Can CLIP-based video learners effectively generalize to video domains they have not encountered during training?
+Building upon the impressive success of CLIP (Contrastive Language-Image Pretraining), recent pioneer works have proposed to adapt the powerful CLIP to video data, leading to efficient and effective video learners for open-vocabulary action recognition.
+Inspired by that humans perform actions in diverse environments, our work delves into an intriguing question: Can CLIP-based video learners effectively generalize to video domains they have not encountered during training?
 To answer this, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, and conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners under various types of domain gaps.
-Our evaluation demonstrates that previous methods exhibit limited action recognition performance in unseen video domains, revealing potential challenges of the cross-domain open-vocabulary action recognition task.
-To address this task, our work focuses on a critical challenge, namely scene bias, and we accordingly contribute a novel scene-aware video-text alignment method.
+The evaluation demonstrates that previous methods exhibit limited action recognition performance in unseen video domains, revealing potential challenges of the cross-domain open-vocabulary action recognition task.
+In this paper, we focus on one critical challenge of the task, namely scene bias, and accordingly contribute a novel scene-aware video-text alignment method.
 Our key idea is to distinguish video representations apart from scene-encoded text representations, aiming to learn scene-agnostic video representations for recognizing actions across domains.
-Extensive experimental results demonstrate the effectiveness of our method.
+Extensive experiments demonstrate the effectiveness of our method.
 The benchmark and code will be available at https://github.com/KunyuLin/XOV-Action/.
 1 Introduction
+Action recognition aims to recognize what actions humans are performing in videos, which has wide applications in surveillance systems, health monitoring, etc [1; 2].
+Recently, inspired by the impressive success of CLIP (Contrastive Language-Image Pretraining) [3] across various image understanding tasks, pioneer works propose to adapt CLIP to video data for action recognition [4; 5; 6; 7; 8; 9].
+Different from traditional models that focus on closed-set recognition [10; 11; 12; 13; 14], this new CLIP-based recognition paradigm leads to efficient video learners with remarkable open-vocabulary recognition abilities, i.e., they achieve state-of-the-art performance for both closed-set and open-set action categories on various video datasets with moderate training cost.
+Such open-vocabulary abilities significantly enhance the practical value of action recognition models.
+As shown in Figure 1 (a), in this work, we delve into an intriguing question: Can CLIP-based video learners effectively generalize to video domains that they have not encountered during training?
+This generalization ability in unseen video domains is crucial for action recognition models in practice, since models often suffer from environment changes when deployed in real-world applications [15; 16].
+For example, surveillance systems will encounter actions performed under illumination shifts caused by day-night change or weather change.
+Therefore, from a practical perspective, we expect that a CLIP-based video learner possessing open-vocabulary abilities can robustly adapt to domain changes for action recognition.
 Figure 1: (a) Can CLIP-based video learners effectively generalize to unseen test domains?
 For example, given a model trained with normal videos, we wonder if it can effectively recognize actions in dark videos.
-(b) We conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners on four test datasets, namely UCF [59], HMDB [30], ARID [78] and NEC-Dr [12].
+(b) We conduct a comprehensive evaluation for five state-of-the-art CLIP-based video learners on four test datasets, namely UCF [17], HMDB [18], ARID [19] and NEC-Dr [20].
 For each test dataset, we report the accuracy of closed-set and open-set action categories according to the training categories in Kinetics400 [13].
 As shown above, these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains.
-Note that, for each metric, we report the best performance among all methods (namely different checkpoints), please refer to Table 3 for the full results.
-Best viewed in color.
-Action recognition aims to recognize what actions humans are performing in videos, which has wide applications in surveillance systems, health monitoring, sport analysis, etc [28, 62, 14, 13, 94, 17, 22].
-Recently, inspired by the impressive success of Contrastive Language-Image Pretraining (CLIP) [52] across various image understanding tasks, pioneer works propose to adapt CLIP to video data for action recognition [72, 46, 26, 75, 53, 73].
-Different from traditional models that focus on closed-set recognition [56, 71, 38, 6, 3], the new CLIP-based recognition paradigm leads to efficient video learners with remarkable open-vocabulary recognition abilities, i.e., they achieve state-of-the-art performance for both closed-set and open-set action categories on various popular video datasets with moderate training cost.
-Such open-vocabulary abilities significantly enhance the practical relevance of action recognition models.
-As shown in Figure 1(a), in this work, we delve into an intriguing question: Can CLIP-based video learners effectively generalize to video domains that they have not encountered during training?
-The generalization ability in unseen video domains is crucial for action recognition models in practice, since models often suffer from environment changes when deployed in real-world applications [7, 85].
-For example, surveillance systems will encounter actions performed under illumination shifts caused by day-night change or weather change.
-Therefore, from a practical perspective, we expect that a CLIP-based video learner possessing open-vocabulary abilities can robustly adapt to domain changes for action recognition.
-To answer this, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, where the four test datasets exhibit various levels of domain gaps in comparison to the training datasets.
-Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners for cross-domain open-vocabulary action recognition.
+Note that, for each metric, we report the best performance among all methods, please refer to Table 3 for the full results.
+To answer this question, we establish the first CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, which includes four test datasets exhibiting various levels of domain gaps in comparison to the training datasets.
+Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners in cross-domain open-vocabulary action recognition.
 As shown in Figure 1 (b), our evaluation reveals that these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains.
@@
 Also, we note that the open-set recognition performance of these CLIP-based video learners is far from reaching saturation even in domains with moderate domain gap (i.e., ACC of 42.22% on HMDB).
-Overall, cross-domain open-vocabulary action recognition is challenging and underexplored.
-To address this task, our work focuses on tackling the domain gaps by mitigating scene bias, which is a critical challenge of the task.
-Specifically, the scene bias is caused by the high correlation between human actions and specific scenes in training videos, also known as the spurious correlation [36, 11, 69, 33].
-For example, as shown in Figure 1(a), humans usually perform jumping in track and field in the training domain (Kinetics400), thus models prefer to recognize the action according to the track and field since the static scenes are easier to fit [36, 11, 69].
-However, humans may perform jumping in hallways (ARID) or basketball courts (NEC-Dr), thus recognizing “jump” by the track and field would result in recognition errors across domains.
+Overall, cross-domain open-vocabulary action recognition poses challenges for existing CLIP-based video learners, as it requires recognizing both closed-set and open-set action categories in unseen video domains.
+To address this task, our work focuses on tackling domain gaps by defending against the scene bias when fitting training videos.
+Specifically, the scene bias is caused by the strong associations between actions and specific scenes in training videos (aka., spurious correlation [21; 22; 23; 24; 25; 26]), and it would hinder models’ generalizability across domains.
+For example, as shown in Figure 1 (a), if humans usually perform jumping on track-and-field grounds in the training domain (Kinetics400), models will prefer to recognize the action based on the track-and-field grounds since static scenes are easier to fit [21; 22; 24].
+However, humans may perform jumping in hallways (ARID) or on basketball courts (NEC-Dr), thus recognizing “jump” based on track-and-field grounds would result in recognition errors across domains.
 To mitigate the scene bias, we propose a novel Scene-Aware video-Text Alignment (SATA) method for learning scene-agnostic video representations.
-Specifically, we propose to distinguish video representations apart from scene-encoded text representations, which encourages the video encoder to downweight the attention on scene information and thus pay more attention to action information.
+Our key idea is to distinguish video representations apart from scene-encoded text representations, which encourages the video encoder to downweight the attention on scene information and thus pay more attention to action information.
+Experiments on XOV-Action demonstrate that our SATA can effectively improve the closed-set action recognition performance across domains while generally maintaining the open-set performance.
 In summary, our contributions are listed as follows:
-  • We establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, including four test domains exhibiting various levels of domain gap with respect to the training domains.
+(1) We establish the first CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, which includes four test domains exhibiting various levels of domain gap in comparison to the training domains.
 We identify closed-set and open-set categories for each test domain, thus providing a comprehensive way to evaluate models across various situations.
-  • We conduct a thorough evaluation of five state-of-the-art CLIP-based video learners based on XOV-Action.
+(2) We conduct a thorough evaluation of five state-of-the-art CLIP-based video learners based on XOV-Action.
 Our experimental analysis demonstrates the limited performance of these methods in cross-domain open-vocabulary action recognition, revealing potential challenges of this task.
-  • We propose a novel scene-aware video-text alignment method, which learns scene-agnostic video representations to recognize actions across domains.
-And, extensive experiments demonstrate the effectiveness of our method.
+(3) We propose a novel Scene-Aware video-Text Alignment (SATA) method, which aims to learn scene-agnostic video representations to recognize actions across domains.
+Extensive experiments demonstrate the effectiveness of our method in tackling domain gaps.
 2 Related Works
-2.1 Action Recognition
 Action Recognition aims to recognize human actions in videos.
-In the past decade, motivated by the success of deep learning [29, 57, 64, 21, 63, 10, 68], many video classification architectures have been proposed.
-Among these architectures, there are three most popular types, namely 2D CNNs, 3D CNNs and Video Transformers.
-Typically, 2D CNNs adopt 2D convolution for spatial modeling, and conduct temporal modeling beyond spatial modeling [71, 93] or embed temporal shift into spatial modeling [38, 55, 60].
-Different from 2D models, 3D CNNs extends 2D convolution to 3D convolution for adapting video data [65, 6, 66, 67, 16, 90, 35].
-By adopting attention mechanisms, Video Transformers expand the receptive field of 2D and 3D CNNs, leading to remarkable performance [19, 45, 89, 91, 3, 2].
+In the past decade, motivated by the success of deep learning [27; 28; 29; 30; 31; 32; 33], many video classification architectures have been proposed.
+These architectures can primarily be categorized into 2D CNNs, 3D CNNs and Video Transformers.
+Typically, 2D CNNs adopt 2D convolution for spatial modeling, and conduct temporal modeling beyond spatial modeling [11; 34] or embed temporal shift into spatial modeling [12; 35; 36].
+3D CNNs extends 2D convolution to 3D convolution for adapting video data [37; 13; 38; 39; 40; 41; 42].
+By adopting attention mechanisms, Video Transformers expand the receptive field of 2D and 3D CNNs, leading to remarkable performance [43; 44; 45; 46; 14; 47].
 Although above models show promising performance for closed-set action recognition, they usually lack the ability to recognize open-set categories.
-2.2 Cross-Domain Action Recognition
+Recently, inspired by the success of CLIP (Contrastive Language-Image Pretraining) [3; 48], some pioneer works propose to adapt CLIP to video data for action recognition [49; 4; 5; 49; 6; 7; 8; 9; 50; 51; 52; 53; 54; 55].
+Owing to the image-text alignment power of CLIP, these CLIP-based video learners are endowed with remarkable open-vocabulary action recognition abilities with moderate training cost.
+These methods adapt CLIP to video data in various ways.
+For example, ActionCLIP [4] stacks temporal fusion layers on top of the image encoder for modeling temporal dynamics, Ju et al.
+[6] cooperate continuous prompting with temporal Transformer, and X-CLIP [5] proposes cross-frame communication attention for temporal modeling.
+OpenVCLIP [9] and FROSTER [54] propose to improve open-set recognition by harnessing the power of raw CLIP.
+In our work, we focus on the underexplored cross-domain setting for open-vocabulary action recognition, which aims to develop generalizable open-vocabulary action recognition models for unseen video domains.
 Cross-Domain Action Recognition aims to learning video classification models by transferring knowledge from source domains to target domains.
-This area includes two tasks, namely domain-adaptive and domain-generalizable action recognition.
-In domain-adaptive action recognition [61, 77, 25, 7, 79, 5, 80], unlabeled videos from target domains are accessible for training, and thus related works usually focusing on developing models oriented to specific target domains.
-Typical works address domain-adaptive action recognition by learning cross-domain invariance, i.e., exploring the shared action cues between the source and target domains [7, 48, 48, 81, 42, 54].
-Other works usually focus on leveraging the multi-modal nature of video data [44, 58, 27, 84, 92].
-Different from domain-adaptive action recognition, domain-generalizable action recognition [85, 50, 39, 51] aims to learn generalizable models in unseen test domains, i.e., videos of the target domains are not accessible during training.
+This area mainly includes two tasks, namely domain-adaptive and domain-generalizable action recognition.
+In domain-adaptive action recognition [56; 57; 58; 15; 59; 60; 61], unlabeled videos from target domains are accessible for training, thus prevailing works usually focusing on developing models oriented to specific target domains.
+Typical works address domain-adaptive action recognition by learning cross-domain invariance [15; 62; 62; 63; 64; 65].
+Other works usually explore to leverage the multi-modal nature of video data [66; 67; 68; 69; 70].
+Differently, domain-generalizable action recognition [16; 71; 72; 73] aims to learn generalizable models in unseen test domains, i.e., videos of the target domains are not accessible during training.
 In the absence of target videos, existing works usually assume specific types of invariance across domains for generalizable action representation learning, e.g., Yao et al.
-assume that local temporal features are more invariant across domains compared with global temporal features [85].
-Instead of invariance mining, Lin et al.
-propose to learn diverse action features for generalization in unseen target domains [39].
+assume that local features are more invariant across domains compared with global features [16].
 Our cross-domain open-vocabulary action recognition task is closely related to domain-generalizable action recognition, since test domains are unseen during training.
-However, our task is much more challenging as we strive for open-vocabulary abilities, leading to significantly different technical designs.
-2.3 Vision-Language Pretraining
-Vision-language pretraining [52, 34, 37, 1, 88, 86, 15, 83, 18] has made great progress in recent years, and one of the most impressive works is Contrastive Language-Image Pretraining (CLIP) [52].
+However, our task is much more challenging as we strive for cross-domain open-vocabulary abilities, leading to significantly different technical designs.
+Vision-Language Pretraining [3; 48; 74; 75; 76; 77; 78] has made great progress in recent years, and one of the most impressive works is CLIP [3].
 By utilizing web-scale paired image-text data for training, CLIP shows robust zero-shot object recognition abilities.
-Also, some advanced works have demonstrated that CLIP can effectively solve specific downstream tasks by efficient adaptation [97, 87, 49, 96, 53, 41].
-Moreover, integrating CLIP with specialized techniques shows remarkable open-vocabulary abilities on various image understanding tasks, such as object detection and segmentation [52, 74, 20, 31, 40].
-Although large-scale image-text pretraining has achieved great success, video-text pretraining still has room for development [76, 32, 70, 43, 23, 9, 53].
+Also, some advanced works have demonstrated that CLIP can effectively solve specific downstream tasks by efficient adaptation [79; 80; 81; 82; 8; 83].
+Moreover, integrating CLIP with specialized techniques shows remarkable open-vocabulary abilities on various image understanding tasks, e.g., object detection and segmentation [84; 85; 86; 49].
+Although large-scale image-text pretraining has achieved great success, video-text pretraining still has room for development [87; 88; 89; 90; 91; 92; 93; 94; 95].
 It is because videos are inherently more complex than images, and large-scale paired video-text datasets are less available.
-Therefore, it is valuable to develop efficient methods for adapting pretrained image-text models to video understanding tasks.
+Therefore, it is valuable to develop methods for adapting pretrained image-text models to video understanding tasks.
 3 The Proposed XOV-Action Benchmark
@@
 This task aims to develop open-vocabulary action recognition models that are generalizable in unseen target domains by training in the source domain, i.e., recognizing both closed-set and open-set action categories in new test domains.
-This task, closely aligning with real-world applications, requires action recognition models to exhibit open-vocabulary classification abilities and demonstrate robustness in handling environmental changes.
-To evaluate models on this task, we establish a CROSS-domain Open-Vocabulary Action recognition benchmark named XOV-Action, which provides a comprehensive way to analyze models across various situations.
+To evaluate models in this task, we establish XOV-Action, a CROSS-domain Open-Vocabulary Action recognition benchmark, which provides a comprehensive way to analyze models across various situations.
 In what follows, we introduce the components of our XOV-Action benchmark in detail, as well as evaluation metrics and potential challenges.
-Benchmark Components: Our proposed XOV-Action benchmark consists of two source datasets for training and four target datasets for test.
-Specifically, the two source datasets used for training are as follows:
+Table 1: Category statistics of four test domains in our XOV-Action. The closed-set and open-set categories are identified according to Kinetics400.
+Test Domains UCF HMDB ARID NEC-Dr
+# of closed-set actions 50 33 6 7
+# of open-set actions 51 18 5 9
+# of all actions 101 51 11 16
+Domain gap  Small Moderate Large Large
+Benchmark Components: Our proposed XOV-Action benchmark consists of two source datasets for training and four target datasets for test. The two source datasets for training are as follows:
 (1) Kinetics400 [13]: One of the most widely-used action recognition datasets, consisting of 400 action categories.
 Videos in Kinetics400 are collected from YouTube, which are usually recorded in normal environments (e.g., normal illumination and weather).
-Existing works of CLIP-based video learners typically use Kinetics400 for training, and we use the original train-validation split following them.
+Existing CLIP-based video learners typically use Kinetics400 for training, and we use the original split following them.
 (2) Kinetics150: A subset of Kinetics400, composed of 150 action categories selected from the full Kinetics400.
-These 150 categories include all the closed-set categories of the four target datasets in comparison to Kinetics400, and the remaining categories are random sampled.
+These 150 categories include all the closed-set categories of the four target datasets in comparison to Kinetics400, and the remaining categories are randomly sampled.
 We will illustrate the definition of closed-set categories later.
-We construct this subset for a detailed analysis to cross-domain open-vocabulary action recognition models.
+We construct this subset to conduct detailed analysis for cross-domain open-vocabulary action recognition models.
 The four target datasets for test are as follows:
@@
 Specifically, Videos in HMDB are collected mainly from movies, and remaining videos are from Prelinger archive, YouTube or Google videos.
-Overall, HMDB has a moderate domain gap compared with the Kinetics domain [80, 82].
+Overall, HMDB has a moderate domain gap compared with the Kinetics domain [61; 96].
 (3) ARID [19]: A dataset consisting of 11 categories of action videos, which are recorded under dark environments.
 These actions include singular person actions (e.g., jump, run) and actions associated with objects (e.g., drink, pick).
-Due to the significant different illumination conditions, ARID exhibits a large domain gap compared with the Kinetics domain (see Figure 1 for an example).
-Developing models that are generalizable in ARID is challenging, when training data are limited to videos with normal illumination.
+Due to the significantly different illumination conditions, ARID exhibits a large domain gap compared with the Kinetics domain (see Figure 1 for an example).
+When training data are limited to videos with normal illumination, developing models that are generalizable in ARID is challenging.
 (4) NEC-Dr [20]: A dataset consisting of 16 categories of action videos, which are recorded by drones in the same basketball court.
@@
 Then, we evaluate models on the four target datasets, which exhibit various levels of domain gap in comparison to the training datasets.
-Please refer to the Appendix for more details of our XOV-Action.
-Evaluation Metrics: First, we illustrate the definitions of closed-set and open-set categories for test domains: (a) The closed-set categories refer to the categories that share similar meanings and have a common lexicon with the categories in the training domain, following a similar approach in previous zero-shot action recognition works [4, 8, 95].
+Please refer to the Appendix for more details of our XOV-Action benchmark.
+Evaluation Metrics: First, we illustrate the definitions of closed-set and open-set categories for test domains: (a) The closed-set categories refer to the categories that share similar meanings and have a common lexicon with the categories in the training domain, following a similar approach in previous zero-shot action recognition works [97; 98].
 (b) The open-set categories refer to the remaining categories that are not involved in the training domain.
@@
 Category statistics are summarized in Table 1.
-In our experiments, we adopt three evaluation metrics:
-(1) The closed-set accuracy measures the recognition performance of closed-set categories, which primarily evaluates the model abilities of tackling domain gaps.
+In our experiments, we adopt three evaluation metrics: (1) The closed-set accuracy measures the recognition performance of closed-set categories, which primarily evaluates the model abilities of tackling domain gaps when fitting training videos.
 (2) The open-set accuracy measures the performance of open-set categories, which evaluates the generalization abilities across both video domains and action categories.
@@
 Note that, some previous works of CLIP-based video learners treat all the categories in UCF and HMDB as open-set for models trained on Kinetics400.
-However, UCF and HMDB have many overlapping categories with Kinetics400, thus these works make an inaccurate assessment of the model’s open-vocabulary abilities.
+However, UCF and HMDB have many overlapping categories with Kinetics400, thus these works make an inaccurate assessment of models’ open-vocabulary abilities.
 Unlike these works, we provide a more accurate way for evaluation by distinguishing between closed-set and open-set categories.
-By evaluating across various domain gaps and different action categories, our proposed XOV-Action benchmark enables us to perform a wide range of analysis for cross-domain open-vocabulary action recognition models.
+By evaluating across various domains and different action categories, our proposed XOV-Action benchmark enables us to perform a wide range of analysis for cross-domain open-vocabulary action recognition models.
 Potential Challenges: Although cross-domain open-vocabulary action recognition is valuable and practical, it is challenging and underexplored.
-Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners, as shown in Table 3.
-The experimental results demonstrate that these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains.
-Specifically, the results reveal three fundamental challenges of the cross-domain open-vocabulary action recognition task as follows:
-(1) How to tackle the domain gap between the training and test domains?
+Based on XOV-Action, we conduct a comprehensive evaluation of five state-of-the-art CLIP-based video learners.
+As shown in Table 3, the experimental results on XOV-Action demonstrate that these CLIP-based video learners exhibit limited performance when recognizing actions in unseen test domains.
+In addition, the results reveal three fundamental challenges of the cross-domain open-vocabulary action recognition task as follows:
+I.
+How to improve the generalization of closed-set categories in unseen test domains?
 Previous CLIP-based video learners exhibit very limited performance in test domains with large domain gaps, i.e., the best closed-set ACCs are only 53.89% and 31.48% on ARID and NEC-Dr, respectively.
-(2) How to improve the generalization abilities for open-set categories in the presence of domain gaps?
+To address this challenge, it is crucial to learn domain-invariant video representations from training domains.
+Our method mainly targets this challenge by focusing on the mitigation of scene bias.
+II.
+How to improve the generalization of open-set categories in the presence of domain gaps?
 The cross-domain open-set recognition performance is far from reaching saturation even in test domains with moderate domain gaps, e.g., the best open-set ACC is only 42.22% on HMDB.
-(3) How to develop a single model that achieves superior performance on all the test domains?
-It is very challenging since existing models demonstrate varied performance across various test domains as shown in Table 3.
-Table 1: Category statistics of four test domains. Closed-set and open-set categories are identified according to Kinetics400.
-Test Domains UCF HMDB ARID NEC-Dr
-# of closed-set actions 50 33 6 7
-# of open-set actions 51 18 5 9
-# of all actions 101 51 11 16
-Domain gap Small Moderate Large Large
+To address this challenge, it is crucial to harness the power of language models for recognizing open-set categories across domains.
+Additionally, another possible direction is to alleviate overfitting of closed-set categories in training domains [9; 54].
+III.
+How to develop a single model that performs excellently for all categories in all test domains?
+This is very challenging, since existing models show varied performance across various domains, and empirically improvements in cross-domain closed-set performance do not guarantee improvements for open-set categories and could even cause a decrease.
+Given the unique characteristics of each test domain and category type, it is crucial to find a trade-off solution, especially when no additional annotations/data are introduced and the model size remains consistent.
 4 Methodology
-In this work, we focus on tackling the domain gaps for cross-domain open-vocabulary action recognition.
-Specifically, we propose to address a critical challenge of the task, namely scene bias.
-To this end, we propose a novel Scene-Aware video-Text Alignment (SATA) method.
-In what follows, we illustrate our method in detail.
-Figure 2: An overview of our proposed Scene-Aware video-Text Alignment (SATA) method.
-Our method basically includes a video encoder and a text encoder for video and text representation extraction, with a contrastive loss for video-text alignment.
-Based on the scene-encoded text prompts, we propose the Scene-Aware Discrimination and Scene-Aware Consistency losses, aiming to learn scene-agnostic video representations for cross-domain open-vocabulary action recognition.
-Best viewed in color.
 4.1 Problem Formulation
-In cross-domain open-vocabulary action recognition, a set of labeled videos $\mathcal{D}=\{(x,y)\}$ from a source domain are given for training, where $x$ and $y$ denote a source video and its ground-truth label index.
+In cross-domain open-vocabulary action recognition, a set of labeled videos $\mathcal{D}=\{(x,y)\}$ from a source domain are given for training, where $x$ and $y$ denote a source video and its ground-truth action label index.
 The source domain consists of $K$ action categories, and the action names (label texts) are denoted by $\{a_{1},a_{2},\dots,a_{K}\}$.
 Given only source videos for training, our goal is to develop a model that is generalizable in unseen target domains, where the source and target domains follow different distributions and different label spaces.
-Following previous open-vocabulary action recognition works, we sample $T$ frames from each video as model input during training and test, i.e., $x=\{x_{1},x_{2},\dots,x_{T}\}$.
+Following previous open-vocabulary action recognition works, we sample $T$ frames from each video as model input during training and test.
+Figure 2: An overview of our proposed Scene-Aware video-Text Alignment (SATA) method.
+Our method basically includes a video encoder and a text encoder for representation extraction, with a contrastive loss for video-text alignment.
+Based on the scene-encoded text prompts, we propose the Scene-Aware Discrimination and Action-Aware Discrimination losses, aiming to learn scene-agnostic video representations for cross-domain open-vocabulary action recognition.
+Best viewed in color.
 4.2 Model Overview
@@
 Following OpenVCLIP [9], we construct a video encoder by integrating patches from neighboring frames in each self-attention layer of the original CLIP image encoder.
-For video $x$, the video representation $e_{x}$ is obtained by the average of local video representations, i.e., $e_{x}=f_{\text{vid}}(x)=1/T\sum_{t=1}^{T}f^{t}_{\text{vid}}(x)$, where each local representation corresponds to one frame.
+For the video $x$, the video representation $e_{x}=f_{\text{vid}}(x)$ is obtained by the average of local video representations (each local representation corresponds to one frame).
 We use the original CLIP text encoder and keep it frozen during training.
-Following previous works [73, 53], we use a video-text alignment loss to adapt video data, which is formulated as follows:
+Following previous works [9; 8], we use a video-text alignment loss to adapt video data, which is formulated as follows:
 $L_{\text{vta}}=-\log{\frac{\exp{(s(e_{x},e_{a_{y}})/\tau)}}{\sum_{k=1}^{K}\exp{(s(e_{x},e_{a_{k}})/\tau)}}},$ (1)
@@
 The function $g(\cdot)$ transforms an action name into a text prompt in the form of “a video of a person [doing something].”, e.g., “a video of a person abseiling.” for the action “abseiling”.
-To mitigate scene bias, we contribute a Scene-Aware Discrimination loss and a Scene-Aware Consistency loss for learning scene-agnostic video representations, which will be illustrated in the next subsection.
-An overview of our proposed SATA is shown in Figure 2.
+To mitigate scene bias, we contribute a Scene-Aware Discrimination loss and a Action-Aware Discrimination loss for learning scene-agnostic video representations, which will be illustrated in the next subsection.
+An overview of our proposed method is shown in Figure 2.
 4.3 Scene-Aware Video-Text Alignment
@@
 Based on these suffixes, we construct scene-encoded text prompts of ground-truth action category for each video, which is in the form of “a video of a person [doing something] [at/on/in the/a scene].” For example, for the action “abseiling”, we obtain $N$ scene-encoded text prompts, e.g., “a video of a person abseiling in the park.”.
-Table 2: Evaluation of CLIP-based video learners trained on Kinetics150 on our XOV-Action benchmark.
-“C”, “O” and “A” denote the accuracy (%) of closed-set, open-set and all action categories, respectively.
-“AVG” denotes the average performance over the four test domains.
+Then, based on the scene-encoded text prompts, we design a Scene-Aware Discrimination loss, which is formulated as follows:
+$L_{\text{scene}}=-\log{\frac{\exp{(s(e_{x},e_{a_{y}}))}}{\exp{(s(e_{x},e_{a_{y}}))}+\sum_{n=1}^{N}\exp{(s(e_{x},\tilde{e}^{n}_{a_{y}}))}}}.$ (2)
+In this loss, $\tilde{e}^{n}_{a_{y}}=f_{\text{txt}}(\tilde{g}(a_{y},n))$ is the representation of a scene-encoded text prompt, which encodes the semantic information of the $n$-th scene.
+The function $\tilde{g}(\cdot,n)$ transforms an action name into a scene-encoded text prompt using the $n$-th scene suffix.
+According to Eq.
+(2), our Scene-Aware Discrimination loss pushes the video representations away from the scene-encoded text representations in video-text alignment.
+In this way, the loss leverages the strong power of CLIP text encoder, which encourages the video encoder to pay less attention to scene information and thus pay more attention to action information, thereby mitigating the scene bias when fitting training videos.
+Action-Aware Discrimination: In our experiments, we find that although the Scene-Aware Discrimination loss $L_{\text{scene}}$ effectively improves the cross-domain action recognition performance for the closed-set categories, it reduces the cross-domain performance for open-set categories.
+Thus, it leads to limited improvement in the overall cross-domain open-vocabulary performance.
+A critical reason is that $L_{\text{scene}}$ pushes a video away from a scene-encoded text prompt in representation space, and this may result in higher similarity between the video and some non-ground-truth action texts in representation space, relative to the scene-encoded text prompt.
+However, since the scene-encoded text prompt encodes semantic information of the ground-truth category (in addition to a scene), the video should have a more different representation from the non-ground-truth action texts, compared with the scene-encoded text prompt.
+For example, consider a video that depicts a person playing basketball in a court.
+This video should have a dissimilar representation to the text prompts of other categories (e.g., “a video of a person kicking soccer.”) compared with the scene-encoded text prompts (e.g., “a video of a person playing basketball in the park.”).
+Overall, this issue would cause some confusion in video representation space.
+Accordingly, to alleviate the degradation of cross-domain open-set performance, we propose a Action-Aware Discrimination loss to constrain video representation learning.
+Relative to the scene-encoded text representations that encodes ground-truth action semantics, our Action-Aware Discrimination loss pushes the video representations away from the text presentations of non-ground-truth categories, which is formulated as follows:
+$L_{\text{action}}=\frac{1}{N(K-1)}\sum\nolimits_{n=1}^{N}\sum\nolimits_{k\not=y}^{K}\max{\left(0,\delta-s(e_{x},\tilde{e}^{n}_{a_{y}})+s(e_{x},e_{a_{k}})\right)},\vspace{-0.1cm}$ (3)
+where $\delta$ is the margin.
+In principle, $L_{\text{action}}$ introduces a constraint for non-ground-truth action text $a_{k}$ ($k\not=y$), i.e., $s(e_{x},\tilde{e}^{n}_{a_{y}})-s(e_{x},e_{a_{k}})\geq\delta$.
+By using both $L_{\text{scene}}$ and $L_{\text{action}}$, our SATA aims to learn a more reasonable video representation space, i.e., $s(e_{x},e_{a_{y}})>s(e_{x},\tilde{e}^{n}_{a_{y}})>s(e_{x},e_{a_{k}})$ when $k\not=y$.
+In summary, the overall training loss of our SATA is given as follows:
+$L=L_{\text{vta}}+\lambda_{\text{scene}}L_{\text{scene}}+\lambda_{\text{action}}L_{\text{action}},$ (4)
+where $\lambda_{\text{scene}}$ and $\lambda_{\text{action}}$ are coefficients for trade-off. During test, we use text presentations of action categories in target domains for classification.
+Table 2: Evaluation of CLIP-based video learners trained on Kinetics150 on XOV-Action.
+“C”, “O” and “A” denote the ACC (%) for closed-set, open-set and all action categories, respectively.
+“AVG” denotes the average ACC over four test domains.
 * indicates the model variant with frozen text encoder.
-The bold and underlined numbers indicate the best and second best.
+The bold/underlined numbers indicate the best/second best.
+We prioritize the average overall ACC as the primary metric in model comparison.
 Models UCF HMDB ARID NEC-Dr AVG
@@
 OpenVCLIP [9] 82.98 63.17 73.03 56.48 44.96 52.41 48.16 27.13 39.73 19.68 13.38 16.45 51.82 37.16 45.40
-SATA (Ours) 85.53 62.01 73.71 58.70 44.78 53.79 47.35 33.54 41.81 27.55 10.97 19.04 54.78 37.82 47.09
-Then, based on the scene-encoded text prompts, we design a Scene-Aware Discrimination loss, which is given as follows:
-$L_{\text{dis}}=-\log{\frac{\exp{(s(e_{x},e_{a_{y}}))}}{\exp{(s(e_{x},e_{a_{y}}))}+\sum_{n=1}^{N}\exp{(s(e_{x},\tilde{e}^{n}_{a_{y}}))}}}.$ (2)
-In this loss, $\tilde{e}^{n}_{a_{y}}=f_{\text{txt}}(\tilde{g}(a_{y},n))$ is the representation of a scene-encoded text prompt, which encodes semantic information of both the ground-truth action $a_{y}$ and a random scene.
-The function $\tilde{g}(\cdot,n)$ transforms an action name into a scene-encoded text prompt using the $n$-th scene suffix.
-Accordingly, our Scene-Aware Discrimination loss pushes the video representations away from the scene-encoded text representations in video-text alignment.
-In this way, the loss leverages the strong power of CLIP text encoder, which encourages the video encoder to pay less attention on scene information and thus pay more attention to action information, thereby bridging the domain gap.
-Scene-Aware Consistency: In addition, we propose a Scene-Aware Consistency loss to constrain the video representation learning.
-Our intuition is that a video should have a more consistent representation with its scene-encoded text prompts compared with text representations of other action categories, since the scene-encoded text prompts encodes semantic information of the ground-truth action category (in addition to scenes).
-For example, consider a training video that depicts a person playing basketball in a court.
-The video should have a similar representation to the scene-encoded text prompts like “a video of a person playing basketball in the park.”, compared with text prompts of other action categories, e.g., “a video of a person playing football.”.
-Specifically, our proposed Scene-Aware Consistency loss is formulated as follows:
-$L_{\text{con}}=$ $-\frac{1}{N+1}\log{\frac{\exp{(s(\hat{e}_{x},e_{a_{y}}))}}{Z}}$ (3)
-$-\frac{1}{N+1}\sum\nolimits_{n=1}^{N}\log{\frac{\exp{(s(\hat{e}_{x},\tilde{e}^{n}_{a_{y}})))}}{Z}},$
-where $\hat{e}_{x}=f_{\text{proj}}(f_{\text{vid}}(x))$ is the video representation after projection $f_{\text{proj}}(\cdot)$, and $Z=\sum_{k=1}^{K}\exp{(s(\hat{e}_{x},e_{a_{k}}))}+\sum_{n=1}^{N}\exp{(s(\hat{e}_{x},\tilde{e}^{n}_{a_{y}}))}$.
-Note that, the projection function is crucial to harmonize our Scene-Aware Discrimination and Scene-Aware Consistency losses, as shown in our experimental analysis.
-Furthermore, we propose a Text-Adaptive Aggregation module to enhance video representations for classification.
-Specifically, we aggregate local video representations by text-adaptive weights instead of average pooling.
-The aggregation module is formulated as $e^{\text{txt}}_{x}=\sum_{t=1}^{T}w^{t}_{x}e^{t}_{x}$, where $e^{t}_{x}=f^{t}_{\text{vid}}(x)$ is the $t$-th local representation.
-And, the weight $w^{t}_{x}$ is as follows:
-$w^{t}_{x}=\frac{1}{K}\sum_{k=1}^{K}\frac{\exp{(s(e^{t}_{x},e_{a_{k}})/\eta)}}{\sum_{m=1}^{T}\exp{(s(e^{m}_{x},e_{a_{k}})/\eta)}},$ (4)
-where $\eta$ is the temperature.
-In this way, we obtain integrated video representations adaptive to specific action categories.
-In our implementation, we use $e^{\text{txt}}_{x}$ instead of $e_{x}$ in Eq.
-(1) and Eq.
-(2) for training, and also use $e^{\text{txt}}_{x}$ for test.
-The overall training loss of our SATA is given as follows:
-$L=L_{\text{vta}}+\lambda_{\text{dis}}L_{\text{dis}}+\lambda_{\text{con}}L_{\text{con}},$ (5)
-where $\lambda_{\text{dis}}$ and $\lambda_{\text{con}}$ are coefficients for trade-off. During test, we use text presentations of action categories in target domains for classification.
-Table 3: Evaluation of CLIP-based video learners trained on Kinetics400 on our XOV-Action benchmark.
+SATA (Ours) 84.89 62.86 73.82 58.40 44.59 53.53 48.78 29.57 41.08 24.54 12.72 18.48 54.15 37.44 46.73
+5 Experiments
+5.1 Experimental Setups
+Implementation Details: For each video, our model takes $T=16$ frames of size $224\times 224$ as inputs.
+Following ViFiCLIP [8], we adopt temporal jitter, multi-scale random spatial crop and color jitter for augmentation during training.
+During test, we use one temporal clip composed of center-cropped frames for inference.
+Regarding the architecture of the video encoder, we adopt ViT-B/32 and ViT-B/16 initialized by CLIP image encoder for Kinetics150 and Kinetics400, respectively.
+For temporal modeling, we set the temporal receptive field of self-attention layers to 7 in the video encoder.
+We ask GPT-4 [99] to output 300 scene suffixes that are common in daily life, and randomly sample $N=50$ scene suffixes for our losses in each batch.
+The parameters are optimized using AdamW optimizer with a batch size of 512, a learning rate of 8e-6, a momentum of 0.9 and a weight decay of 1e-3.
+If not specified, we set the loss coefficients as $\lambda_{\text{scene}}=0.2$ and $\lambda_{\text{action}}=0.2$, the margin as $\delta=0.5$, and the temperature coefficient as $\tau=0.01$.
+Following OpenVCLIP [9], we adopt the SWA technique [100] for producing the final model.
+All experiments are conducted by PyTorch with 8 NVIDIA RTX 3090 GPUs.
+Table 3: Evaluation of CLIP-based video learners trained on Kinetics400 on XOV-Action.
 We use the officially released checkpoints of previous methods for evaluation, and we use one temporal clip during inference for a fair comparison.
-“C”, “O” and “A” denote the ACC (%) of closed-set, open-set and all action categories, respectively.
-“AVG” denotes the average performance over the four test domains.
+“C”, “O” and “A” denote the ACC (%) for closed-set, open-set and all action categories, respectively.
+“AVG” denotes the average ACC over four test domains.
 * indicates the model variant with frozen text encoder.
-The bold and underlined numbers indicate the best and second best.
+The bold/underlined numbers indicate the best/second best.
+We prioritize the average overall ACC as the primary metric in model comparison.
 Models UCF HMDB ARID NEC-Dr AVG
@@
 OpenVCLIP [9] 89.80 77.28 83.51 58.08 42.22 52.48 45.64 26.36 37.91 8.29 10.28 9.32 50.45 39.04 45.81
-SATA (Ours) 88.51 65.17 76.78 61.44 44.03 55.29 54.49 28.05 43.89 28.70 11.84 20.06 58.29 37.27 49.01
+SATA (Ours) 88.51 65.70 77.05 61.54 41.98 54.63 56.94 25.61 44.38 32.41 10.53 21.18 59.45 35.95 49.31
 Table 4: Ablation study of our proposed SATA trained on Kinetics150.
-“C”, “O” and “A” denotes the accuracy (%) of closed-set, open-set and all action categories, respectively.
-“AVG” denotes the average performance over the four test domains.
-The closed-set ACC primarily evaluates the model abilities of tackling domain gaps.
+“C”, “O” and “A” denotes the ACC (%) for closed-set, open-set and all action categories, respectively.
+“AVG” denotes the average ACC over four test domains.
+The closed-set accuracy primarily evaluates the model abilities of tackling domain gaps when fitting training videos.
 Models UCF HMDB ARID NEC-Dr AVG
@@
 Baseline 84.04 62.22 73.08 57.59 44.59 53.00 47.35 26.52 39.00 21.30 14.04 17.57 52.57 36.84 45.66
-+$L_{\text{dis}}$ 84.63 60.43 72.47 59.41 43.10 53.65 49.18 25.31 39.61 25.93 10.53 18.03 54.79 34.84 45.94
-+$L_{\text{dis}}$&$L_{\text{con}}$ 84.84 62.28 73.50 58.30 44.03 53.26 49.39 29.57 41.44 26.39 11.40 18.70 54.73 36.82 46.73
-Full 85.53 62.01 73.71 58.70 44.78 53.79 47.35 33.54 41.81 27.55 10.97 19.04 54.78 37.82 47.09
-+$L_{\text{dis}}$&$L_{\text{con}}$ w/o Proj 84.20 61.75 72.92 58.50 42.91 53.00 48.78 28.35 40.59 20.14 14.04 17.01 52.91 36.76 45.88
-5 Experiments
-5.1 Experimental Setups
-Implementation details: For each video, our model takes $T=16$ frames of size $224\times 224$ as inputs.
-Following ViFiCLIP [53], we adopt temporal jitter, multi-scale random spatial crop and color jitter for augmentation during training.
-During test, we use one temporal clip composed of center-cropped frames for inference.
-Regarding the architecture of the video encoder, we adopt ViT-B/32 and ViT-B/16 initialized by CLIP image encoder for Kinetics150 and Kinetics400, respectively.
-For temporal modeling, we set the temporal receptive field of self-attention layers to 7 in the video encoder.
-We ask GPT-4 [47] to output 300 scene suffixes that are common in daily life, and randomly sample $N=50$ scene suffixes for Eq.
-(2) in each batch.
-The projection function $f_{\text{proj}}(\cdot)$ is an MLP with residual.
-The parameters are optimized using AdamW optimizer with a batch size of 512, a learning rate of 8e-6, a momentum of 0.9 and a weight decay of 1e-3.
-By default, we set the loss coefficients as $\lambda_{\text{dis}}=0.2$ and $\lambda_{\text{con}}=0.2$, and the temperature coefficients as $\tau=0.01$ and $\eta=0.1$.
-Following OpenVCLIP [73], we adopt the SWA technique [24] for producing the final model.
++$L_{\text{scene}}$ 84.63 60.43 72.47 59.41 43.10 53.65 49.18 25.31 39.61 25.93 10.53 18.03 54.79 34.84 45.94
++$L_{\text{scene}}$&$L_{\text{action}}$ 84.89 62.86 73.82 58.40 44.59 53.53 48.78 29.57 41.08 24.54 12.72 18.48 54.15 37.44 46.73
 Evaluation Protocol: We adopt five state-of-the-art CLIP-based video learners for evaluation and comparison.
@@
 During inference, we use one temporal clip for a fair comparison.
-We use our proposed XOV-Action benchmark for evaluation, which includes test domains of various types of domain gaps and action categories of different types.
-We report the accuracy of closed-set, open-set and all categories in individual test domains as well as the average accuracies, for a comprehensive analysis.
-When conducting a comparison between models, we prioritize the average ovarall accuracy as the primary metric.
+We evaluate each checkpoint on all the four test domains of our XOV-Action benchmark, where different test domains have various types of domain gaps and differ in action category.
+For a comprehensive analysis, we report the accuracy of closed-set, open-set and all categories in individual test domains as well as the average accuracies.
+When conducting a comparison between different methods, we prioritize the average overall accuracy as the primary metric.
 5.2 Results
@@
 These results reveal potential challenges of the task, as discussed in Section 1 and 3.
-By learning scene-agnostic video representations, our proposed SATA outperforms previous state-of-the-arts by 1.69% and 1.45% in terms of average overall accuracy for Kinetics150 and Kinetics400, respectively.
-These improvements are significant considering that the metric represents an average across four accuracy metrics.
-Figure 3: Quantitative analysis of the total number of scene suffixes.
-Figure 3: Quantitative analysis of the total number of scene suffixes.
-Figure 4: Quantitative analysis of the coefficient $\lambda_{\text{dis}}$ for Scene-Aware Discrimination loss by closed-set accuracy on four test domains.
-The horizontal axis shows the value of $\lambda_{\text{dis}}$.
+By learning scene-agnostic video representations, our proposed SATA outperforms previous state-of-the-arts by 1.33% and 1.75% in terms of average overall accuracy for Kinetics150 and Kinetics400, respectively.
+These improvements are significant, especially considering that the metric is an average across four test domains.
 Ablation Study: Table 4 summarizes the results of our ablation study.
 In the table, the baseline refers to the model trained with the basic video-text alignment loss $L_{\text{vta}}$.
-By introducing our proposed Scene-Aware Discrimination loss $L_{\text{dis}}$, our model obtains improvement on XOV-Action in terms of closed-set ACC on all the four test domains.
-Specifically, our model obtains significant improvements of 1.83% and 4.63% in terms of closed-set ACC on ARID and NEC-Dr, which have large domain gaps with the training domain.
-This is because that our proposed Scene-Aware Discrimination loss encourages the video encoder to downweight the attention on scene information in videos and thus pay more attention to action information that are generalizable across different video domains.
-However, we find that introducing only the Scene-Aware Discrimination loss reduces the open-set performance.
-By using both Scene-Aware Discrimination loss $L_{\text{dis}}$ and Scene-Aware Consistency loss $L_{\text{con}}$, our model can obtain improved closed-set performance while maintaining the open-set performance, and this leads to an improvement of 1.07% in terms of average overall accuracy over the baseline.
-Notably, the projection within our Scene-Aware Consistency loss is crucial to harmonize the two losses.
-If the projection is dropped, the improvement is very limited as shown in the last line of Table 4.
-Finally, by introducing the Text-Adaptive Aggregation module, our model obtains further improvement in terms of average overall accuracy (please refer to the Appendix for more analysis of this module).
+By introducing our proposed Scene-Aware Discrimination loss $L_{\text{scene}}$, our model obtains improvement on XOV-Action in terms of closed-set accuracy on all the four test domains.
+Specifically, our model obtains significant improvements of 1.83% and 4.63% in terms of closed-set accuracy on ARID and NEC-Dr, which have large domain gaps with the training domain.
+This is because that $L_{\text{scene}}$ encourages the video encoder to downweight the attention on scene information in videos and thus pay more attention to action information that are generalizable across different domains.
+However, we find that introducing only the Scene-Aware Discrimination loss reduces the open-set performance, leading to limited improvement in average overall accuracy on XOV-Action.
+By further introducing the Action-Aware Discrimination loss $L_{\text{action}}$, our model can obtain improved cross-domain closed-set accuracy over the baseline while maintaining the open-set accuracy on each test domain as much as possible, and this leads to an improvement of 1.07% in terms of average overall accuracy over the baseline.
+Figure 3: Quantitative analysis of the total number of scene suffixes.
+Figure 4: Quantitative analysis of the coefficient $\lambda_{\text{scene}}$ for the Scene-Aware Discrimination loss by closed-set accuracy on four test domains.
+The horizontal axis shows the value of $\lambda_{\text{scene}}$.
+Effect of Scene-Aware Discrimination: We conduct a quantitative analysis to the loss coefficient $\lambda_{\text{scene}}$, and the results on the four test domains are shown in Figure 4.
+The results in the figure show that our model consistently obtains better closed-set performance on all test domains with a larger loss weight.
+The comprehensive results demonstrate that our proposed Scene-Aware Discrimination loss effectively tackles various types of domain gaps by mitigating scene bias in fitting training videos.
 Effect of Scene Suffixes: In Figure 4, we show how the suffix number affects the model performance.
-We set the loss coefficient $\lambda_{\text{dis}}=0.5$ in this experiment.
+To show the loss effect more clearly, we set the loss coefficient $\lambda_{\text{scene}}=0.5$ in this experiment.
 As shown in the figure, our model obtains better closed-set accuracy as the total number of scene suffixes increases.
-According to the design of our Scene-Aware Discrimination loss, we distinguish videos apart from more scene-encoded text prompts in feature space with more scene suffixes used.
+According to the design of our Scene-Aware Discrimination loss, we distinguish
+Figure 5: Qualitative analysis by t-SNE [101].
+videos apart from more scene-encoded text prompts in representation space with more scene suffixes used.
 In this way, the video representations are less likely to be confused with scenes, leading to stronger abilities for recognizing actions in unseen domains.
-Figure 5: Analysis of the coefficient $\lambda_{\text{con}}$ for Scene-Aware Consistency loss. The horizontal axis shows the value of $\lambda_{\text{con}}$.
-Effect of the Scene-Aware Discrimination loss: We conduct a quantitative analysis of the loss coefficient $\lambda_{\text{dis}}$, and the results on the four test domains are shown in Figure 4.
-The results in the figure show that our model consistently obtains better closed-set performance on all test domains with a larger loss weight.
-The comprehensive results demonstrate that our proposed Scene-Aware Discrimination loss effectively bridges the domain gaps of various types by mitigating scene bias.
-Effect of the Scene-Aware Consistency loss: We conduct a quantitative analysis of the loss coefficient $\lambda_{\text{con}}$, and the results are shown in Figure 5.
-As shown in the figure, our model obtains higher open-set performance but lower closed-set performance when using a larger loss weight.
-The results empirically reveal that a good trade-off between our Scene-Aware Discrimination and Consistency losses is crucial for cross-domain open-vocabulary action recognition.
-Figure 6: Qualitative analysis by attention visualization. Best viewed in color.
-Qualitative analysis: In Figure 6, we show a qualitative analysis by attention visualization.
-As shown in the figure, compared with the baseline that pays much attention to scenes, our model focuses on the action performers when recognizing actions across domains.
-The results demonstrate that our model effectively mitigates scene bias.
-5.3 Conclusion
+Effect of Action-Aware Discrimination: Figure 5 intuitively shows how the Action-Aware Discrimination loss $L_{\text{action}}$ affects representation space.
+As shown in the figure, without the $L_{\text{action}}$ introduced, our Scene-Aware Discrimination loss may lead to higher similarity between a video and some non-ground-truth action texts compared with the scene-encoded text prompts in representation space, which would lead to confusion in video classification.
+By introducing $L_{\text{action}}$, our SATA produces a much more reasonable representation space, where a video is more distant from the non-ground-truth action texts compared with the scene-encoded text prompts.
+For the analysis to the loss weight $\lambda_{\text{action}}$, please refer to the Appendix.
+6 Conclusion
 This work concentrates on a valuable but underexplored task, namely cross-domain open-vocabulary action recognition.
-By contributing the XOV-Action benchmark, we provided a comprehensive way to evaluate models across various types of domain gaps and different action categories, and our evaluation of state-of-the-art CLIP-based video learners revealed the challenging nature of this valuable task.
-To address this task, we proposed a novel scene-aware video-text alignment method, aiming to tackle the domain gap by learning scene-agnostic video representations.
-Extensive experimental results demonstrated the effective of our method.
-The cross-domain open-vocabulary action recognition task presents numerous challenges that remain to be addressed, such as cross-category generalization in cross-domain settings.
-Despite our efforts, we focus only on resolving scene bias in our current methodology.
-However, our proposed benchmark provides a comprehensive way for evaluation.
+By contributing the XOV-Action benchmark, we provided a comprehensive way to evaluate models across various types of domain gaps and different action categories, and our evaluation of five state-of-the-art CLIP-based video learners revealed the challenging nature of this valuable task.
+To address this task, we proposed a novel Scene-Aware video-Text Alignment (SATA) method, aiming to mitigate the scene bias when fitting training videos.
+Quantitative comparison on XOV-Action showed that our SATA can effectively improve the closed-set action recognition performance across domains while generally maintaining the open-set performance, and extensive analysis demonstrated the effectiveness of our method.
+This work focuses on resolving the scene bias in current methodology, which mainly targets the cross-domain generalization for closed-set categories (i.e., Challenge I discussed in Section 3).
+Despite our efforts, the task of cross-domain open-vocabulary action recognition still presents numerous challenges, including improving open-set recognition across domains and developing a unified model for all cases (Challenges II and III in Section 3).
+Fortunately, our proposed XOV-Action benchmark provides a comprehensive way for evaluation and analysis.
 We believe that our work will serve as a catalyst for further advancement, and we hope it will inspire future innovative solutions for this field.